How many grams of BCl3 contains 0.83 x 1023 molecules of BCl3

1
Expert's answer
2021-03-23T05:54:34-0400

Molar Mass of BCl3 = 117.17 g/mol

6.022 × 10^23 = 117.17 g

1 molecule = 117/(6.022 × 10^23)

0.83× 10^ 23= (117.17 ×0.83× 10^23) /(6.022×10^23)

(117.17 × 0.83)/ 6.022 = 16.1493g